OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

jimthompson5910 (jim_thompson5910):

use y-y1 = m(x-x1) m = slope (x1,y1) = point the line goes through

OpenStudy (anonymous):

A ???

jimthompson5910 (jim_thompson5910):

Similar Example Write the point-slope form of the equation passing through (3, -2) with a slope of 9. y-y1 = m(x-x1) y-y1 = 9(x-x1) ... plug in the given slope y-(-2) = 9(x-3) ... plug in the given point y+2 = 9(x-3) This isn't your answer since it's just an example

OpenStudy (anonymous):

So C

jimthompson5910 (jim_thompson5910):

good